a pharmacist has 20% solution of phenol and a 60% phenol solution. how much of each solution should be mixed together to form 800 milliliters of a 47.5% solution?
Answer by checkley77(12844)

.60x+.20(800-x)=.475*800
.60x+160-.20x=380
.40x=380-160
.40x=220
x=220/.40
x=550 ml. of 60 solution is used.
800-550=250 ml. of 20% solution is used.
Proof:
.60*550+.20*250=380
330+50=380
380=380